OpenStudy (anonymous):

could someone please explain how to do this step by step? \[\sqrt{x-7}-2=0\]

OpenStudy (anonymous):

add 2 to both sides to get \[\sqrt{x-7}=2\] square both sides to get \[x-7=4\] add 7 to both sides to get \[x=11\]

OpenStudy (anonymous):

check is \[\sqrt{11-7}-2=\sqrt{4}-2=2-2=0\]

OpenStudy (lgbasallote):

*sigh* i have been waiting for ages to point out a mistake of @satellite73 when will that ever happen =_=

OpenStudy (anonymous):

satellite73 are you going to be on for a bit? I may need more help on problems? And HUGE TY TY TY

OpenStudy (lgbasallote):

satellite is a busy guy :p i would love to help you...but i ahve finals in a few minutes

OpenStudy (anonymous):

i turn in to a pumpkin at the stroke of 9:30

OpenStudy (anonymous):

go ahead and post, happy to help @lgbasallote i make mistakes constantly, but i prefer to call them "typos"

OpenStudy (anonymous):

that is what this is for, I need to write detailed notes on how to do these and other problems, he gave example questions Sigh LOL

OpenStudy (anonymous):

LOL ok, I wish now I was a better note taker bigtime, or that my teacher was a better explainer LOL